[Brahms-l] Shift report 20040115 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 15 2004 - 16:14:23 EST
------------------------------------------------------------
Date: Thu Jan 15 16:14:18 2004, Shift: 08:00-16:00
Supervisor: catalin, with: 
Log book#15 pages 33 - 33

Shift Summary:
------------------------------------------------------------
1. Quite long store with low ZDC rates from 11 am until 1530.
2. Runs 9689, 9692, 9696, 9697 finishing the 4degs settings at B1219 magnetic field.
3. Quite a lot of problems during this run: tiles hv is down, there are problems with T5. Dana worked on getting the vertex distribution narrower.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 15 16:17:53 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 15 2004 - 16:18:08 EST